OpenStudy (anonymous):

MEDAL + FAN if f(X)=|2x-3|+|x+1| find f`(2) , f`(-2), f`(0)

OpenStudy (anonymous):

Abs(2x-3)=sqrt((2x-3)^2)

OpenStudy (anonymous):

Same with the other term

OpenStudy (anonymous):

abs(x+1)=sqrt((x+1)^2)

OpenStudy (anonymous):

so just take the derivative with the new terms and you're golden

OpenStudy (anonymous):

the answers are:4,8,4

OpenStudy (whpalmer4):

@AnnieDJ204 no, the answers are not 4,8,4. You have misread the problem as asking you to evaluate \(f(x)\) rather than \(f'(x)\) which is the first derivative of \(f(x)\). Completely different animals. This problem asks you to find the slope of the function \(f(x) = |2x-3|+|x+1|\) at \(x = 2,\, x = -2,\,\text{and } x = 0\) The attached graph shows \(|2x-3|, |x+1|, \text{ and } |2x-3|+|x+1|\) None of the line segments have a slope of 4 or 8 at any point, as you can see.

OpenStudy (anonymous):

thank u for correcting it.